# Learning an Ensemble Token from Task-driven Priors in Facial Analysis

### v1

@@ 5 Limitations and future works
The proposed ET-Fuser framework exhibited performance gains across tasks and evaluation metrics, while maintaining a demonstrably minimal computational overhead, with particularly notable improvements observed at the feature representation level. However, a salient limitation inherent in ensemble token methodologies is their susceptibility to overfitting datasets. The training phrase, which relies on task priors and presents each encoder with caonnical input, potentially impedes the learning of multi-modal or multi-input modalities. We posit that this research serves as a cornerstone towards the effective exploitation of task priors and encourage future investigations to broaden its applicability beyond the domain of facial analysis, extending to more generalized domains.

### v2

@@ 5 Conclusion
The proposed KT-Adapter framework exhibited performance gains across tasks and evaluation metrics, while maintaining a demonstrably minimal computational overhead, with particularly notable improvements observed at the feature representation level. However, a salient limitation inherent in knowledge token methodologies is their susceptibility to overfitting datasets. The training phase, which relies on task priors and presents each encoder with canonical input, potentially impedes the learning of multi-modal or multi-input modalities. We posit that this research serves as a cornerstone towards the effective exploitation of task priors and encourage future investigations to broaden its applicability beyond the domain of facial analysis, extending to more generalized domains.